# Quantifying the Impact of Motion on 2D Gaze Estimation in Real-World Mobile Interactions

### v1

@@ Abstract
Keywords Mobile Eye Tracking  $\cdot$ 2D Gaze Estimation  $\cdot$ Calibration  $\cdot$ Mobile Devices  $\cdot$ User Studies
@@ 3.1 Eye Tracking Module
The server is equipped with the PyTorch and scikit-learn frameworks [44, 45] to decouple gaze estimation models that cannot run at high speeds on mobile devices for long periods of time due to battery limitations. The model is based on the  [33] architecture, pre-trained on GazeCapture [33] and fine-tuned on device-specific data from the RGBDGaze dataset [4]. The model has an average Euclidean error of 2.05 cm on the test set and can be used as a reasonable feature extractor for our 2D gaze estimation task.
@@ 3.4 Motion Sensing Module
Since no out-of-the-box API measures head distance, we developed a custom method based on the linear proportional relationship between interpupillary distance and the distance from the eyes to the phone screen. We manually measured the physical interpupillary distance ($IOD_{\text{Physical}}$) and the distance from the interpupillary centre to the camera ($D_{\text{Head-Camera}}$) using a ruler. The interpupillary distance within the image frame ($IOD_{\text{Frame}}$) is then calculated based on facial landmarks detected in the images. These measurements are used to establish a regression model, as shown in Eq. 1, to learn the coefficients $\beta_{0}$, $\beta_{1}$, and the random error $\epsilon$, allowing us to approximate $D_{\text{Head-Camera}}$. This method is implemented in the motion sensing module to provide real-time estimations of head-to-camera distance, as depicted in Figure 2.
@@
Figure 2: Head-to-camera distance calculation.
@@ 6.2.2 Impact of Mobility on 2D Gaze Estimation
First we conduct an in-depth analysis by comparing the gaze estimation performance on two settings: (1) base model performance; that is, directly inputting the facial landmarks and facial images to the iTracker model in Section 3.1 and obtaining the gaze points; (2) one-off calibration approach; that is, training one calibrator using any of the calibration data following the procedure in Section 3.2 and then using this calibrator for gaze point prediction. This references the default procedure of commercial eye-trackers [51, 17], where a calibration is performed once before using the eye-tracker, and then the same calibrator is applied until the entire experiment is completed.
@@ 7.1.2 Clustering distribution and user behaviour
7.2 Limitations of 2D daze estimation in handling dynamic head poses

### v2

@@ Abstract
Keywords Mobile Gaze Tracking  $\cdot$ 2D Gaze Estimation  $\cdot$ Calibration  $\cdot$ Mobile Devices  $\cdot$ IMU Sensors  $\cdot$ User Studies
@@ 3.1 Eye Tracking Module
The server is equipped with the PyTorch and scikit-learn frameworks [44, 45] to decouple gaze estimation models that cannot run at high speeds on mobile devices for long periods of time due to battery limitations. The model is based on the iTracker [33] architecture. It was initially pre-trained on the GazeCapture dataset [33] and achieved an average Euclidean error of 2.05 cm on its test set. And subsequently fine-tuned using device-specific data from the RGBDGaze dataset [4], achieved an average Euclidean error of 1.76 cm on its test set, and can be used as a reasonable feature extractor for our 2D gaze estimation task. In our studies, this fine-tuned iTracker-based model serves two roles: (1) it provides direct gaze predictions, referred to as the "base model" performance, and (2) it acts as a feature extractor, where its penultimate layer’s activations are used to train participant-specific Support Vector Regression (SVR) calibrators as detailed in Section 3.2.
@@ 3.4 Motion Sensing Module
Since no out-of-the-box API measures head distance, we developed a custom method based on the linear proportional relationship between interpupillary distance and the distance from the eyes to the phone screen. We manually measured the physical interpupillary distance ($IOD_{\text{Physical}}$, typically in millimeters or centimeters) and the distance from the interpupillary centre to the camera ($D_{\text{Head-Camera}}$) using a ruler. The interpupillary distance within the image frame ($IOD_{\text{Frame}}$) is then calculated based on facial landmarks detected in the images. These measurements are used to establish a regression model, as shown in Eq. 1, to learn the coefficients $\beta_{0}$, $\beta_{1}$, and the random error $\epsilon$, allowing us to approximate $D_{\text{Head-Camera}}$. This method is implemented in the motion sensing module to provide real-time estimations of head-to-camera distance, as depicted in Figure 2. While this custom regression-based estimation provides a practical means of approximating head-to-camera distance, it is acknowledged that it may introduce a minor source of measurement error compared to dedicated depth sensors; however, initial calibration of this method using known distances aimed to minimize this.
@@
Figure 2: Head-to-camera distance calculation using interpupillary distance (IPD) regression.
@@ 6.2.2 Impact of Mobility on 2D Gaze Estimation
First, we conduct an in-depth analysis by comparing gaze estimation performance across several configurations:
  1. 1.
    Base model performance: This refers to the direct gaze predictions obtained from our fine-tuned iTracker-based model (as described in Section 3.1) when applied to the User Study 2 dataset without any participant-specific calibration data from this particular study. This serves as our uncalibrated baseline for this study.
  2. 2.
    One-off calibration approach: This involves training an SVR calibrator (as per Section 3.2) using features extracted by the fine-tuned iTracker-based model from a single set of calibration data collected during User Study 2. This SVR calibrator then adjusts the subsequent gaze predictions.
  3. 3.
    Other participant-specific calibration approaches (task-specific, motion-specific): These also utilize features from the fine-tuned iTracker-based model to train SVR calibrators, but with varying subsets of User Study 2 calibration data as detailed later in this section.
The one-off calibration approach references the default procedure of commercial eye-trackers [51, 17], where a calibration is performed once before using the eye-tracker, and then the same calibrator is applied until the entire experiment is completed.
@@ 7.1.2 Clustering distribution and user behaviour
7.2 Limitations of 2D gaze estimation in handling dynamic head poses